OpenStudy (anonymous):

I need help with operation signs involving polynomial simplification. Here is always where I fail in algebra. I especially get confused when simplifying terms that are several operation signs apart from each other. Some help, please? Here's an example. {n - 1 - [n - 1 - (n - 1)]}

OpenStudy (whpalmer4):

Be systematic and careful :-) \[{n - 1 - [n - 1 - (n - 1)]} = {n - 1 - [n - 1 - n + 1]}\] because - -1 = +1 \[ {n - 1 - [n - 1 - n + 1]} = {n - 1 - n -(-1) - (-n) - 1} = \]\[n - 1 - n + 1 + n - 1 = n -1\]